OpenStudy (anonymous):

What is the limit of 4-x as x approaches 3? Would it simply be 1? If so, can someone explain why it's 1?

OpenStudy (anonymous):

vs. it doesn't exist

Parth (parthkohli):

Yes, it'd be 3.

hartnn (hartnn):

x approaches 3 means x is very very near to 3 but not exactly 3,say 2.999999 or 3.0000001...to find the limiting value,we just put x=the number it approaches....so here it would be 4-3=1

Parth (parthkohli):

The function \(4 - x\) is continuous, so we can just plug in what \(x\) is approaching.

Parth (parthkohli):

Here, \(x \to 3\), but we can plug-and-play since the function is continuous.\[x \to 3, f(x ) \to 4 - 3=3 \]

Parth (parthkohli):

4 - 3 = 1*

OpenStudy (anonymous):

Alright I understand, thank you very much!